OpenStudy (anonymous):

Write an equation using this: Passes through (1,2) with slope -5/3.

OpenStudy (noseboy908):

use point-slope form: y-y1=m(x-x1)

OpenStudy (anonymous):

use y=mx+b x=1 y=2 m=-5/3 solve for b. 2=(-5/3)(1) + b 2+(5/3)=b b=3.6666 therefore your equation is: y=(-5/3)x + 3.66